If 10mg of benzene is subjects to combustion analysis, what mass of CO2 and H2O will be produced?
1
Expert's answer
2020-10-16T08:24:45-0400

balanced equation

2C6H6       (g)+ 15 O2      (g)                 12 CO2 (g ) +      6 H2O  (g)

moles of benzene

2(12*6+ 1*6)=156

1mole=156g

   ? = 10g

there for

(1*10/156)=0.0641


Mole ratio

2:15:12:6


moles of carbondioxide

  moles= ( 0.0641*12/2)=0.3846 mole

mass=moles *molar mass

       0.3844*44= 16.9136g


moles of H20

0.0641*6/2= 0.1923moles

mass=moles *moler mass

          0.1923*18

        =3.4614 g